• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name brads

    Log entry time 14:01:07 on March 19, 2012

    Entry number 367173

    keyword=gep/g2p update -- shifts canceled through Tuesday OWL; Tuesday DAY on standby

    Damage assessment of the septum magnet is in progress. The techs were able to remove a damaged electrical junction this morning and will be running some low-power tests on the coil packs themselves this afternoon.

    We expect to have a clearer picture of what happened, and what options are available going forward, in time for the 1:30pm Hall A meeting tomorrow (Tuesday).

    For now we will cancel all Monday SWING, and all Tuesday OWL shifts.

    Tuesday DAY TO an 3rd shifts are on standby: please email brads@jlab.org with a phone number at which you could be reached on Tuesday if we need you. We ask that the Tuesday DAY shift leader staff the counting house as a point of contact.

    I will send out another update as soon as possible on Tuesday to clarify shift needs for Tuesday SWING and later.



    A copy of this log entry has been emailed to: g2p